OpenStudy (anonymous):

What are the solutions of 5x2 = 20x?

OpenStudy (anonymous):

\[5x^2 - 20x = 0\] Take 5x common: \[5x(x-4) = 0\] Separately put them = 0: \[x = 0 \qquad and \qquad (x-4) = 0\] So, \[x = 0 \qquad and \qquad x= 4\]

OpenStudy (anonymous):

Thanks